Two trains going in opposite directiosn leave at the same time.Train B travels 5 mph faster than train A. In 5 hours the trains are 425 miles apart.Find the speed of each.




Thanks.

Answer by rfer(16322)

425/5=85 mph total
x+x+5=85
2x=80
x=40 mph
x+5=45 mph